# Breaking the Modality Barrier: Universal Embedding Learning with Multimodal LLMs

### v1

@@ 2.1. Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) extend LLMs to process and integrate cross-modal information (Bin et al., 2024; Wang et al., 2024b, b; Jiang et al., 2024a). A seminal development in this field is LLaVA (Liu et al., 2023), which encodes images into token representations using vision encoders like CLIP (Radford et al., 2021; Sun et al., 2023a) and projects them into the LLM’s token space. Following this breakthrough, numerous MLLM variants (Peng et al., 2023; Lin et al., 2024; Zhang et al., 2024a; Zhu et al., 2023) have demonstrated remarkable performance in multimodal comprehension and reasoning tasks. CogVLM (Wang et al., 2023) incorporates a trainable visual expert module within the attention and feed-forward network layers of the language model, yielding substantial performance enhancements across 17 canonical cross-modal benchmarks. LLaVA-1.6 (Liu et al., 2024c) introduces the “AnyRes” technique to process variable high-resolution images, significantly improving fine-grained visual understanding. Phi3.5-Vision (Abdin et al., 2024), a 4.2-billion parameter model evolved from phi-3.5-mini, exhibits strong reasoning capabilities for both single and multi-image text prompts. While these advancements, the autoregressive next-token prediction objective of MLLM inherently constrains its capacity to learn efficient multimodal representations.
@@ 3.1. Preliminary
Unlike the dual-tower structure of CLIP, MLLM incorporates three essential components: a vision tower, a projection layer, and an LLM backbone. This unified structure supports flexible processing of both unimodal (image or text) and multimodal (interleaved image-text) inputs, enabling diverse task execution within a single framework. In this work, we present a novel two-stage framework that enables MLLMs to learn universal multimodal embedding for diverse downstream tasks. In the first textual discrimination knowledge distillation stage, we follow the previous work (Jiang et al., 2024b) and employ the prompt: ”¡Text¿ Summarize above sentences in one word: n” to guide the LLM to compress textual information into a single embedding at the last token. In the second hard negative enhanced instruction tuning stage, we utilize task-specific prompts from VLM2Vec (Jiang et al., 2025) to adapt UniME fit diverse downstream tasks such as: ”¡Image¿ Represent the given image with the following question: ¡Text¿” for Visual Question Answering (VQA) tasks and ”¡Image¿ Find a caption for the news in the given photo.” for retrieval tasks.
@@ 3.2. Textual Discriminative Knowledge Distillation
Inspired by previous research (Jiang et al., 2024b), we enhance the LLM backbone of the MLLM to improve its overall embedding capabilities. The autoregressive decoder architecture of LLMs, constrained by a causal masking mechanism, inherently restricts their discriminative capacity and poses challenges in effectively distinguishing between diverse items. To address this limitation, we introduce discriminative textual knowledge distillation (as shown in Figure 2), transferring knowledge from the state-of-the-art LLM-based embedding model NV-Embed V2 (Lee et al., 2024), which employs multiple diverse datasets and removes the causal attention mask during contrastive training. Specifically, we first decouple the LLM component from the MLLM architecture and process text-only inputs using the embedding prompt: “Summarize the above sentences in one word: n”. Then we obtain the normalized student text embeddings and teacher text embeddings (which is extracted offline) from the hidden state of the final token, where is the batch size, is the dimension of the embeddings. Subsequently, we implement discriminative distribution alignment by minimizing the Kullback-Leibler (KL) divergence (Shlens, 2014) between the embeddings from the teacher model and student model:
@@ A.1. Training Details
A.2. MMEB Benchmarks
A.3. Retrieval Task Evaluation Benchmarks

### v2

@@ 2.1. Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) extend LLMs to process and integrate cross-modal information (Bin et al., 2024; Wang et al., 2024b, b; Jiang et al., 2024a). A seminal development in this field is LLaVA (Liu et al., 2023), which encodes images into token representations using vision encoders like CLIP (Radford et al., 2021; Sun et al., 2023a; Yang et al., 2023; Gu et al., 2024; Yang et al., 2025; Hu et al., 2025) and projects them into the LLM’s token space. Following this breakthrough, numerous MLLM variants (Peng et al., 2023; Lin et al., 2024; Zhang et al., 2024a; Zhu et al., 2023) have demonstrated remarkable performance in multimodal comprehension and reasoning tasks. CogVLM (Wang et al., 2023) incorporates a trainable visual expert module within the attention and feed-forward network layers of the language model, yielding substantial performance enhancements across 17 canonical cross-modal benchmarks. LLaVA-1.6 (Liu et al., 2024c) introduces the “AnyRes” technique to process variable high-resolution images, significantly improving fine-grained visual understanding. Phi3.5-Vision (Abdin et al., 2024), a 4.2-billion parameter model evolved from phi-3.5-mini, exhibits strong reasoning capabilities for both single and multi-image text prompts. While these advancements, the autoregressive next-token prediction objective of MLLM inherently constrains its capacity to learn efficient multimodal representations.
@@ 3.1. Preliminary
Unlike the dual-tower structure of CLIP, MLLM incorporates three essential components: a vision tower, a projection layer, and an LLM backbone. This unified structure supports flexible processing of both unimodal (image or text) and multimodal (interleaved image-text) inputs, enabling diverse task execution within a single framework. In this work, we present a novel two-stage framework that enables MLLMs to learn universal multimodal embedding for diverse downstream tasks. In the first textual discrimination knowledge distillation stage, we follow the previous work (Jiang et al., 2024b) and employ the prompt: “¡Text¿ Summary above sentences in one word: n” to guide the LLM to compress textual information into a single embedding at the last token. In the second hard negative enhanced instruction tuning stage, we utilize task-specific prompts from VLM2Vec (Jiang et al., 2025) to adapt UniME fit diverse downstream tasks such as: ”¡Image¿ Represent the given image with the following question: ¡Text¿” for Visual Question Answering (VQA) tasks and ”¡Image¿ Find a caption for the news in the given photo.” for retrieval tasks.
@@ 3.2. Textual Discriminative Knowledge Distillation
Inspired by previous research (Jiang et al., 2024b), we enhance the LLM backbone of the MLLM to improve its overall embedding capabilities. The autoregressive decoder architecture of LLMs, constrained by a causal masking mechanism, inherently restricts their discriminative capacity and poses challenges in effectively distinguishing between diverse items. To address this limitation, we introduce discriminative textual knowledge distillation (as shown in Figure 2), transferring knowledge from the state-of-the-art LLM-based embedding model NV-Embed V2 (Lee et al., 2024), which employs multiple diverse datasets and removes the causal attention mask during contrastive training. Specifically, we first decouple the LLM component from the MLLM architecture and process text-only inputs using the embedding prompt: “¡Text¿ Summary above sentences in one word: n”. Then we obtain the normalized student text embeddings and teacher text embeddings (which is extracted offline) from the hidden state of the final token, where is the batch size, is the dimension of the embeddings. Subsequently, we implement discriminative distribution alignment by minimizing the Kullback-Leibler (KL) divergence (Shlens, 2014) between the embeddings from the teacher model and student model:
@@ A.1. Training Details
A.2. Retrieval Task Evaluation Benchmarks